# InstructAny2Pix: Flexible Visual Editing via Multimodal Instruction Following

### v2

@@ Abstract
The ability to provide fine-grained control for generating and editing visual imagery has profound implications for computer vision and its applications. Previous works have explored extending controllability in two directions: instruction tuning with text-based prompts and multi-modal conditioning. However, these works make one or more unnatural assumptions on the number and/or type of modality inputs used to express controllability. We propose InstructAny2Pix, a flexible multi-modal instruction-following system that enables users to edit an input image using instructions involving audio, images, and text. InstructAny2Pix  consists of three building blocks that facilitate this capability: a multi-modal encoder that encodes different modalities such as images and audio into a unified latent space, a diffusion model that learns to decode representations in this latent space into images, and a multi-modal LLM that can understand instructions involving multiple images and audio pieces and generate a conditional embedding of the desired output, which can be used by the diffusion decoder. Additionally, to facilitate training efficiency and improve generation quality, we include an additional refinement prior module that enhances the visual quality of LLM outputs. These designs are critical to the performance of our system. We demonstrate that our system can perform a series of novel instruction-guided editing tasks. The code is available at https://github.com/jacklishufan/InstructAny2Pix.git
@@
1 Introduction
With growing fidelity of generative models of images [3, 22], the ability to control their outputs is critical for many real-world applications, ranging from creative generation to synthetic augmentations. However, popular state-of-the-art tools such as ControlNet[35] and T2I[21] can only perform specific edits they are trained on. Many of these models also require inputs such as Canny edge, Depth, or Surface Normal, making them inaccessible to general users without expertise in computer vision. To mitigate this, instruction-based image editing methods such as InstructPix2Pix[4] allow users to describe their instructions in natural languages, such as “add a dog.” But, such methods are still limited to simple instruction on which they are trained and cannot generalize to complex instructions involving multiple editing operations or multiple objects. Additionally, they cannot take additional audiovisual inputs, making them struggle with tasks such as style transfer.
We propose InstructAny2Pix, the first instruction-following image editing system that can follow complicated, multi-modal, multi-object instructions. Concrete examples of such instructions can be “add the [sound] to [image],” where the sound can be that of a dog barking or a piece of music. It can also be “add [object A] and remove [object B] from [image],” where objects can be represented by either images, text, or audio. Additionally, it can also include free-form instructions like “change [image A] to the style of [image B]” or “fit [image] to [music].” Our work represents a significant expansion in the scope of image editing instructions. Through both quantitative and qualitative evaluations, we show that our proposed method achieves high performance on diverse editing tasks.
@@ 1 Introduction
First, we train our model on large amounts of multi-modal edit instructions. Our training data consists of diverse instructions, including adding audio to an image, adding an image to an image, adding multiple objects, or replacing an object with another object where objects are represented by either images, text, or audio. We generate this dataset by prompting state-of-the-art Large Language Models with human-written examples for each category of instructions. This diverse dataset is crucial for extending instruction-following capabilities of image-editing systems to long, complex, and multi-modal instructions.
Second, we integrate recent techniques in instruction tuning of language models and multi-modal representation learning to efficiently understand instructions involving inputs of multiple modalities. In particular, we encode the multi-modal inputs using the encoder and insert them into the instruction sequence that is passed to an LLM. The LLM then learns to understand the instruction and generate an output embedding. This embedding is then used to condition a diffusion model that generates the image.
@@ 2 Related Works
The second family is based on text prompts. These methods only require prompts describing the desired output, making them more accessible. The naive approach is to perform DDIM[27] inversion that converts an image back into the latent space of a diffusion model and generates a new image from such a latent representation. Prompt2Prompt (P2P) [10] improves this baseline by injecting the cross-attention maps during the diffusion process. Plug-and-Play [31] additionally injects convolution features to provide more refined control. Parallel to these lines of work, Null-text Inversion [20] proposed a better inversion method for input images through a learnable null text prompts. While these methods are easier to use than the previous family, they still require detailed descriptions of each desired output. This limits their usability in applications like batched editing.
The last family is instruction-guided methods. Unlike previous methods, these models only requires a vague text instruction such as “add fireworks”. InstructPix2Pix [4] achieves this by curating a large machine-generated image editing dataset using P2P, then directly fine-tuning on instruction-following generation. MagicBrush [34] curated a higher quality human-annotated dataset by requesting humans to perform editing operations using tools such as PhotoShop. MGIE [7] utilizes a multi-modal large language model to process editing instructions and input images. But it still operate on a single source image with text-only instructions. All these works are limited to simple instructions and require iteratively applying the model recurrently for complex instructions. Additionally, they cannot take audiovisual input as auxiliary information.
Unlike previous works, our work extends the scope of valid instructions to multi-modal, multi-object inputs and can perform complicated instructions in a single turn.
Figure 3: Examples of generated pairs from the MM-Inst dataset showing many image variations.
@@ 3 Methods
Our pipeline is illustrated in Figure 1. To generate a sample of an edited image given multi-model instructions, we first leverage a pretrained multi-modal encoder that encodes different modalities into a unified embedding space. We then pass the multimodal instruction through a large language model to generate an embedding sequence. We compute input embeddings of the instruction using a large language model (LLM) and substitute the embeddings of the non-linguistic tokens with these embeddings obtained via multi-modal encoder. The large language model eventually generate an output embedding [gen]. Since there can be multiple input images in the instruction, we let the LLM generate an additional embedding [base] used to retrieve the target image to be edited. A diffusion model then generates the edited image using this embedding and retrieved image.
To maximize training efficiency and provide better control over edited outputs, we train the LLM and diffusion model individually instead of end-to-end. During training, the LLM learns to predict the representation of the edited images. The diffusion model learns to generate an image based on the representation of that image in the same embedding space. We also incorporate a refinement prior module refines the output embedding of the LLM based on human preference.
In particular, given an input text sequence $T$ containing reference tokens such as [image], [audio] referring to multi-modal inputs $\{X_{\text{i}}\}_{\text{i=1:n}}$, we first generate the embedding of $\{H_{\text{i}}\}_{\text{i=1:n}}=F_{\text{enc}}(\{X_{\text{i}}\}_{\text{i=1:n}})$ where $F_{\text{enc}}$ is the multi-modal encoder. We then obtain the embedded representation $H_{\text{T}}=E(T)$ where $E$ is the input embedding layer of LLM. The next step is to insert $\{H_{\text{i}}\}_{\text{i=1:n}}$ into the corresponding location of $H_{\text{T}}$. A projection layer $P_{\text{{enc}}}$ is used to project the $\{H_{\text{i}}\}_{\text{i=1:n}}$ to the embedding space of the LLM. This insertion leads to a new sequence of embeddings $H_{\text{T}}^{\prime}$.
We extract the last hidden state $H_{\text{LLM}}=F_{\text{LLM}}(H_{\text{T'}})$ from the LLM after passing $H_{\text{T}}^{\prime}$ as input. We then locate the position corresponding to two special tokens [base] and [gen]. This gives us two embeddings $H_{\text{base}},H_{\text{gen}}$. An additional projector $P_{\text{out}}$ is employed to map the embedding back to the embedding space of the encoder.
We retrieve the base image id $k$ by finding the embedding with the highest similarity with $H_{\text{base}}$. We then perform DDIM inversion on the source image $X_{\text{k}}$ and invert it to a vector $z_{\text{k}}$ in the latent space of the diffusion model through DDIM inverse. The diffusion process is then performed on the latent $z_{\text{k}}$ of the inverted image using condition $H_{\text{gen}}$, which leads to an output image $X_{\text{out}}$ that preserves the spatial structure of the original image while incorporating desired changes. The inference process can be described by Algorithm 1.
Algorithm 1 InstructAny2Pix— Inference
input trained networks $F_{\text{enc}},F_{\text{LLM}},F_{\text{diffusion}}$
${H_{\text{i}}}_{\text{i=1:n}}\leftarrow F_{\text{enc}}(\{X_{\text{i}}\}_{\text{i=1:n}})$ // Encode Multi-Modal:
$H_{\text{T}}\leftarrow E(T)$ // Encode Text
$H_{\text{T'}}\leftarrow\text{insert}(H_{\text{T}},P_{\text{enc}}(\{H_{\text{i}}\}_{\text{i=1:n}}))$ // Substitute
$H_{\text{LLM}}\leftarrow F_{\text{LLM}}(H_{\text{T'}})$
$H_{\text{base}},H_{\text{gen}}\leftarrow P_{\text{out}}(\text{extract}(H_{\text{LLM}}))$ // Extract Output
$k\leftarrow\text{argmax}_{i}(\text{Sim}(H_{i},H_{\text{base}}))$ // Retrieval
$z_{\text{k}}\leftarrow\text{inverse}(X_{\text{k}})$ // DDIM Inversion
$z_{\text{k}}^{\prime}\leftarrow\alpha z_{\text{k}}+(1-\alpha)\epsilon;\epsilon\sim\mathcal{N}(0,\boldsymbol{I})$
$z_{\text{k}}^{\prime}\leftarrow z_{\text{k}}^{\prime}\lVert z_{\text{k}}\rVert/\lVert z_{\text{k}}^{\prime}\rVert$ // Mixing Latent with Noise
$H_{\text{gen}}^{\prime}\leftarrow F_{\text{prior}}(H_{\text{gen}})+H_{\text{gen}}+\beta H_{\text{k}}$ // Mixing Conditions
$X_{\text{out}}\leftarrow F_{\text{diffusion}}(z_{\text{k}},H_{\text{gen}}^{\prime}/\|H_{\text{gen}}^{\prime}\|)$ // Generative Editing
return Edited image $X_{\text{out}}$
Where $F_{\text{prior}}$ is the refinement prior, $\epsilon$ is i.i.d Gaussian noise introduced to add randomness in generation, $\alpha$ controls such randomness, and $\beta$ is a hyper-parameter controlling the strength of edits.
3.1 Multi-Modal Encoder
We use ImageBind [9] as our encoder. ImageBind is trained on multi-modal data, including image, audio, depth, and heat maps. We take the image and audio encoder of ImageBind as our multi-modal encoder. We observe that different modalities have different norm distributions, so we perform $L_{2}$ normalization on ImageBind features before passing them to the LLM. During training, we keep the encoder frozen. The only trainable part is the projection layer.
3.2 Instruction Following Multi-Modal LLM
We use Vicuna-7b [29] as our base language model. The input sequence consists of a text with multiple references to images and audios, such as “remove [audio] from [image]” and “add [image 1] to [image 2].” The input text is projected to a latent space through the LLM’s input embedding layer. This leads to a sequence $H_{\text{T}}\in\mathbb{R}^{L\times D}$ where $L$ is the sequence length and $D$ is the dimension of the latent space. We then identify the locations in this sequence that correspond to references of multi-modal inputs like “[image],” “[audio]” and replace them with corresponding multi-modal features. Since the dimensions of the two embedding spaces are not necessarily identical, we incorporate a MLP projector $P_{\text{enc}}$ prior to the substitution. This gives us $H_{\text{T}}^{\prime}\in\mathbb{R}^{L\times D}$ which is identical to $H_{\text{p}}$ except in locations corresponding to reference words such as “image” and “audio”.
We add two new output tokens to extend the output of the LLM: namely “[base]” and “[gen]”. The [base] token is used to retrieve the desired source image when multiple inputs are present, and the [gen] token is used to produce the conditioning embeddings for the diffusion model. In particular, we apply an MLP projection layer $P_{\text{out}}$ to the hidden states from the last LLM layer and obtain $H_{\text{base}},H_{\text{gen}}$. We compute the cosine similarity of $H_{\text{base}}$ and images features from encoder to retrieve the source image, which is then used together with $H_{\text{gen}}$ to condition the diffusion models.
During training, the model is trained with a cross-entropy loss for auto-regressively predicting output tokens and a regression loss for the $H_{\text{base}},H_{\text{gen}}$ embeddings. In particular, the loss can be formulated as:
$\mathcal{L}_{\text{LLM}}=\mathcal{L}_{\text{ce}}+\lVert H_{\text{base}}-F_{\text{enc}}(X_{\text{base}})\rVert_{2}^{2}+\\ \lVert H_{\text{gen}}-F_{\text{enc}}(Y)\rVert_{2}^{2}$ (1)
where $X_{\text{base}}$ is the target image and $Y$ is the edit outcome.
3.3 Refinement Priors
While it is possible to fully align the output of the LLM to the image embedding space, we discovered that in practice, the LLM fails to converge given a reasonable amount of training time and data because of its large size and slow training. To efficiently train our system, we make use of a Refinement Prior which learns to restore a corrupted image embedding to its original state. We consider two kinds of corruptions: Gaussian noise and domain shift. For Gaussian noise, we simply sample standard Gaussian noise and add it to the ground truth image embedding. For domain shifts, we make use of image-text and image-audio pairs and learn to predict the ImageBind image embedding given corresponding text or audio embedding.
We use a decoder-only transformer as our Refinement Prior. To further enhance the refinement capability, we also use the LAION aesthetic score [26] $f$ of an image as additional input. This facilitates the model to generate samples that are visually pleasing to human eyes. Given an image $X$, the corruption operator $C(.)$ that randomly adds noise to the embedding or retrieves the paired embedding of a different modality, let $F_{\text{ase}}$ be LAION aesthetic predictor, our loss for the prior module is:
$\mathcal{L}_{\text{{prior}}}=\lVert F_{\text{enc}}(X)-F_{\text{prior}}(C\circ F_{\text{enc}}(X),F_{\text{ase}}(X))\rVert_{2}^{2}.$ (2)
We find that this refinement scheme allows us to improve generation quality without jointly training the LLM and the prior, even though the prior is not explicitly learning to refine the LLM output. We hypothesize that our corruption operator is simply general enough.
While this process involves training on large amounts of paired multi-modal data, including image and audio, the feature extraction can be performed offline, and the final extracted feature dataset only takes up less than 70GB of disk space. The training itself can be performed on eight 24GB Nvidia A5000 GPUs in less than 72 hours.
As a side effect, our corruption scheme gives us the ability to use the prior as paired multi-modal generation and perform tasks such as text-to-image and audio-to-image without instructions. We will provide some brief qualitative evaluation on this feature and leave a more rigorous study for future research since this work focuses on image editing.
3.4 Datasets and Training
3.4.1 Reconditioning Diffusion models
The diffusion model is conditioned on the embedding generated by the multi-modal encoder and trained using the standard diffusion loss [11]:
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{t}\left[\lVert\epsilon_{t}-F_{\text{diffusion}}(z_{t},F_{\text{enc}}(X))\rVert_{2}^{2}\right]$ (3)
where $z_{t}$ is a noised version of the latent in the diffusion process and $\epsilon_{t}$ is the noise at time $t$. We use 2M samples of high-quality images from LAION-Aesthetic-V2 datasets [26] and SDXL [22] as our base model.
3.4.2 Prior Alignment

### v3

@@ Abstract
The ability to provide fine-grained control for generating and editing visual imagery has profound implications for computer vision and its applications. Previous works have explored extending controllability in two directions: instruction tuning with text-based prompts and multi-modal conditioning. However, these works make one or more unnatural assumptions on the number and/or type of modality inputs used to express controllability. We propose InstructAny2Pix, a flexible multi-modal instruction-following system that enables users to edit an input image using instructions involving audio, images, and text. InstructAny2Pix  consists of three building blocks that facilitate this capability: a multi-modal encoder that encodes different modalities such as images and audio into a unified latent space, a diffusion model that learns to decode representations in this latent space into images, and a multi-modal LLM that can understand instructions involving multiple images and audio pieces and generate a conditional embedding of the desired output, which can be used by the diffusion decoder. Additionally, to facilitate training efficiency and improve generation quality, we include an additional refinement module. These designs are critical to the performance of our system. We demonstrate that our system can perform a series of novel instruction-guided editing tasks.
Keywords:
Image Editing Multi-Modal Text-to-Image
1 Introduction
@@
With growing fidelity of generative models of images [4, 24], the ability to control their outputs is critical for many real-world applications, ranging from creative generation to synthetic augmentations. However, popular state-of-the-art tools such as ControlNet[37] and T2I[22] can only perform specific edits they are trained on. Many of these models also require inputs such as Canny edge, Depth, or Surface Normal, making them inaccessible to general users without expertise in computer vision. To mitigate this, instruction-based image editing methods such as InstructPix2Pix[5] allow users to describe their instructions in natural languages, such as “add a dog." However, such methods are still limited to simple instruction on which they are trained and cannot generalize to complex instructions involving multiple editing operations or multiple objects. Additionally, their inability to take additional audiovisual inputs, makes them struggle with tasks such as style transfer.
We propose InstructAny2Pix, the first instruction-following image editing system that can follow complicated, multi-modal, multi-object instructions. Concrete examples of such instructions can be “add the [sound] to [image]," where the sound can be that of a dog barking or a piece of music. It can also be “add [object A] and remove [object B] from [image]," where objects can be represented through either images, text, or audio. Additionally, Our system also accommodates free-form instructions like “change [image A] to the style of [image B]" or “fit [image] to [music]." This work significantly broadens the scope of image editing instructions. Through both quantitative and qualitative evaluations, we show that our proposed method achieves high performance on diverse editing tasks.
@@
First, we train our model on large amounts of multi-modal edit instructions. Our training data consists of diverse instructions, including incorporating audio into an image, adding an image to an image, adding multiple objects, or replacing an object with another object where objects are represented by either images, text, or audio. We generate this dataset by prompting state-of-the-art Large Language Models with human-crafted examples for each category of instructions. This diverse dataset is crucial for extending instruction-following capabilities of image-editing systems to long, complex, and multi-modal instructions.
Second, we integrate recent techniques in instruction tuning of language models and multi-modal representation learning to efficiently understand instructions involving inputs of multiple modalities. In particular, we encode the multi-modal inputs using the encoder and insert them into the instruction sequence that is passed to a LLM. The LLM then learns to understand the instruction and generate an output embedding. This embedding is then used to condition a diffusion model that generates the image.
@@ 2.1 Image Editing
The second family is based on text prompts. These methods only require prompts describing the desired output, making them more accessible. The naive approach is to perform DDIM[29] inversion that converts an image back into the latent space of a diffusion model and generates a new image from such a latent representation. Prompt2Prompt (P2P) [11] improves this baseline by injecting the cross-attention maps during the diffusion process. Plug-and-Play [32] additionally injects convolution features to provide more refined control. Parallel to these lines of work, Null-text Inversion [21] proposed a better inversion method for input images through a learnable null text prompts. While these methods are easier to use than the previous family, they still require long detailed descriptions of each desired output. This limits their usability in applications like batched editing.
The last family is instruction-guided methods. Unlike previous methods, these models only requires a vague text instruction such as “add fireworks". InstructPix2Pix [5] achieves this by curating a large machine-generated image editing dataset using P2P, then directly fine-tuning on instruction-following generation. MagicBrush [36] curated a higher quality human-annotated dataset by requesting humans to perform editing operations using tools such as PhotoShop. MGIE [7] utilizes a multi-modal large language model to process editing instructions and input images. But it still operate on a single source image with text-only instructions. All these works are limited to simple instructions and require iteratively applying the model recurrently for complex instructions. Additionally, they cannot take audiovisual input as auxiliary information.
Unlike previous works in this area, our work extends the scope of valid instructions to multi-modal, multi-object inputs and can perform complicated instructions in a single turn.
2.2 Multi-Modal Conditioned Generation
Parallel to these image editing methods, there are some previous attempts to achieve image generation with multi-modal conditioning using Multi-modal language models. BLIP-Diffusion [15] incorporates BLIP [16] as a multi-modal encoder that generates subject embedding for the diffusion model. Using this approach, it was able to generate images that following text prompts and reference images. Kosmos-G [23] also makes use of multi-modal language models. By directly aligning the representation space of multi-modal language models with that of a diffusion model, Kosmos-G allows image generation following multiple reference images. However, since these works focus on generation rather than instruction-based image editing, they requires compositional text inputs. These typically follow patterns like “[x] in [y] wearing [z]". They do not support removal and replacement of object, and general free-form instructions They also cannot faithfully respect the spatial structure of reference images.
Audio guided image generation is a relatively uncharted area. AAI [34] achieves sound-guided generation by aligning audio representation to reference images. This method is very limited in that it requires retrieving 3-5 reference images and performing gradient descent optimization steps for each audio input at inference time.
Unlike previous works in this area, our work is the first to support interleaved audiovisual inputs and free-form image editing instructions.
Figure 3: Data generation Pipeline. 1. We prompt LLM with sampled captions and examples to generate a diverse set of instructions. 2. We generate a candidate dataset using a variety of caption or segmentation based methods, and filter them into a final dataset using CLIP score. 3. We show some example image pairs in the filtered dataset.
@@ 3 Methods
Figure 4: Training pipeline of InstructAny2Pix. 1.Language Alignment 2. Generation Learning 3. Refinement Learning 4. Realigning Diffusion Model 5.Instruction Finetuning
@@
Our pipeline is illustrated in Figure 1. To generate a sample of an edited image given a multi-modal instruction, we first leverage a pretrained multi-modal encoder to encode non-textual inputs into a set of embeddings. We then pass these embeddings alongside the text instruction to a large language model and generate a sequence of control tokens. Each control token belongs to one of the following type: [base], [gen], [sub]. The embedding of the [base] token is used to retrieve the source image of editing operation. The embedding of [sub] tokens is used to retrieve reference images of relevant subjects in the desired output. Typically there is only one [base] token, but there can be more than one [sub] tokens to account for multiple reference images. The embedding of [gen] tokens is used to condition the diffusion model. It will be further refined by the refinement module. Finally, the diffusion model generates an image output from retrieved source image, the retrieved embeddings of reference images and refined embeddings of [gen] tokens.
The training pipeline of InstructAny2Pix  consists of five stages: Multi-Modal Language Alignment, Multi-Modal Generation Learning, Refinement Learning, Realigning Diffusion Models and Instruction Guided Finetuning.
3.1 Multi-Modal Language Alignment
We use a MLP projector to map the output of multi-modal encoder to the embedding space of LLM. The goal of multi-modal alignment stage is to learn a good mapping between these two spaces. We make use of a large number of image-text pairs and audio-text pairs. We freeze all parts of the model other than the projector. We prompt the model with the text instruction “describe the [x]", where [x] is the embedding of an image or audio piece. We use the ground truth caption as the target sequence. Standard cross-entropy loss is applied to the output of the LLM.
3.2 Multi-Modal Generation Learning
We use a MLP projector to map the output of LLM back to the representation space of the multi-modal encoder. For clarity, we refer this projector as the output projector, and the projector described in Sec. 3.1 as the input projector. We freeze all parts of the model other than the input and output projector. We make use of image-text pairs and image-audio pairs in this stage. We prompt the model with instruction “generate [x]", where [x] is either a ground truth caption or an embedding of an audio piece. We use the ground truth image embedding encoded by the multi-modal encoder as the target, and apply L2 loss on the output projector. Since the final linear head of LLM is frozen, we cannot obtain [gen] token in the output. We assume the [gen] sequence comes immediately after user input, and directly retrieve the last-layer embedding of LLM at corresponding location as inputs to the output projector.
3.3 Refinement Learning
We observe that there are many low quality images used in the alignment step, most of which come from audio-image pairs. Particularly, many of these pairs come from internet videos with low resolution, bad exposure, undesired motion blur, and other artifacts. We empirically verified this observation by evaluating the LAION aesthetic scores of the dataset. To mitigate the effect of low quality data, we incorporate a refinement module, which is a transformer that learns to improve the image quality in the embedding space. Since pairs consisting of one image and its improved version is hard to obtain, we make use existing image-text and image-audio pairs. As the embedding space of the multi-modal encoder is aligned, we can directly model $\text{Pr}(x|e,s)$, where $x$ is the image, $e$ is the embedding and $s$ is aesthetic score. We can use the text and audio embedding during the training, and substitute them with image embedding at inference.
In this stage, only the refinement module is used. Given image-text and image-audio pairs, the refinement module takes the aesthetic score of the image and the embedding of text or audio as the input, and learns to predict the embedding of the ground-truth image. We apply the L2 regression loss on the predictions.
3.4 Realigning Diffusion Models
The goal of this stage is to realign a text-to-image diffusion model to accept conditional embeddings in the representation space of multi-modal encoder. In particular, it should learn to generate image from both refined [gen] embeddings which encode some global semantics and reference image embeddings retrieved by [sub] token. We employ an MLP projector to map the embeddings to the representation space of the diffusion model. In this stage, we train U-Net of the diffusion model and the MLP projector, keeping the VAE of the diffusion model unchanged. We only use high quality image filtered by aesthetic score in this stage and train the model in a self-supervised fashion. We use the ground truth image embedding as [gen] input, and image embedding of a cropped region as [sub] input.
There are several challenges: first, there may be a distribution shift between the ground truth image embedding and those generated by the refinement module; second, in the training process the [gen] and [sub] input comes from the ground truth image, hence it is important to prevent the model from overfitting to the global embedding and ignore the region embedding. To address the first challenge, we add random gaussian noise to [gen] embedding to improve the robustness. To address the second challenge, we randomly drop the global and/or region embedding independently at a probability of 0.5. Additionally, we randomly perform unconditioned inpainting in the cropped region so that [gen] embedding comes from an image with the region corresponding to [sub] repainted. Because inpainting is expensive, we pre-select a subset and generate these images offline.
In this stage, we apply the standard diffusion loss as [26]. The regions used to crop the image are either generated by an object detector or randomly sampled at 1:1 ratio. The regions used for inpainting are all generated by detector.
3.5 Instruction Guided Finetuning
In this stage, we fine tune the LLM and projectors end-to-end on multi-modal edit instructions and pairs of source image and edit results. We defer the details of data generation process to Sec. 3.7. The input projector is used to map the encoded embeddings of reference images and audio pieces to the representation space of LLM, while the output projector is used to map the LLM output back to the representation space of multi-modal encoder. We formulate the target output sequence as “[base] [sub] .. [sub] [gen] .. [gen]" where each [sub] corresponding to a reference image. We only add a “[sub]" token if the referenced object should appear in the desired output. For example, in the following instructions “add [a]", “remove [b]", “replace [c] with [d]", only [a] and [d] will have corresponding [sub] tokens in the target sequence. We apply cross entropy classification loss on the output of the final linear head of LLM, as the standard language training. Additionally, we fetch the corresponding embeddings of target sequence from the last hidden layer of LLM and project them into the representation space of multi-modal encoder through output projector. L2 regression loss is applied between the [base] embedding and ground truth embedding of the source image, between [sub] embeddings and ground truth embeddings of reference images, and between the [gen] embedding and ground truth embeddings of the target image. The diffusion model is not used in this process. We do not need to actually generate images as the loss is directly applied in the embedding space.
3.6 Implementation Details
We use ImageBind [10] as the encoder and SDXL[24] as the decoder. We use Vicuna-7B[30] as the base LLM. We use two-layer MLP with GELU activation for the input and output projectors.
3.7 Datasets
3.8 Paired Dataset